# MAPS: Advancing Multi-Modal Reasoning in Expert-Level Physical Science

### v1

@@ Abstract
Pre-trained on extensive text and image corpora, current Multi-Modal Large Language Models (MLLM) have shown strong capabilities in general visual reasoning tasks. However, their performance is still lacking in physical domains that require understanding diagrams with complex physical structures and quantitative analysis based on multi-modal information. To address this, we develop a new framework, named Multi-Modal Scientific ReAsoning with Physics Perception and Simulation (MAPS) based on an MLLM. MAPS decomposes expert-level multi-modal reasoning task into physical diagram understanding via a Physical Perception Model (PPM) and reasoning with physical knowledge via a simulator. The PPM module is obtained by fine-tuning a visual language model using carefully designed synthetic data with paired physical diagrams and corresponding simulation language descriptions. At the inference stage, MAPS integrates the simulation language description of the input diagram provided by PPM and results obtained through a Chain-of-Simulation process with MLLM to derive the underlying rationale and the final answer. Validated using our collected college-level circuit analysis problems, MAPS significantly improves reasoning accuracy of MLLM and outperforms all existing models. The results confirm MAPS offers a promising direction for enhancing multi-modal scientific reasoning ability of MLLMs. We will release our code, model and dataset used for our experiments upon publishing of this paper.
@@ 1 Introduction
To address the aforementioned limitations, we proposed Multi-Modal Scientific ReAsoning with Physics Perception and Simulation (MAPS), a novel framework for solving complex multi-modal reasoning problems in physical disciplines. And in this paper we verified the effectiveness of MAPS in electrical discipline, which typically involves multiple circuit diagrams and is representative of expert-level physical science requiring reasoning on complex physical diagrams. The core idea of MAPS is to decompose expert-level reasoning problems into two sub-tasks: understanding the physical diagram and reasoning based on this comprehension and related physical knowledge.
MAPS realizes physical diagram understanding by fine-tuning a visual language model using carefully designed synthetic data, resulting in what we term the Physics Perception Model (PPM). The role of PPM is to translate physical diagrams into simulation language descriptions that can be executed by a simulator. At the inference stage, MAPS integrates the converted simulation language description and their respective simulation results, obtained through a Chain-of-Simulation process, to derive intermediate rationales and ultimately the final answer to the question. Experiment results on college-level circuit analysis problems demonstrate that our framework can successfully address the challenges in complex multi-modal reasoning tasks in physical science. Most importantly, it significantly reduces the occurrence of hallucination when using and solving physical equations. This advancement creates new avenues for precise multi-modal scientific reasoning using MLLMs.
@@ 2 Motivation
Solution to Issue in Perception: Translate physical diagrams into simulation language descriptions. We adopt simulation language for two reasons. First, it describes the physical scene of the diagram using a formal language, enabling the language model to directly access the fundamental structure behind the question.
Second, with parameters of physical objects provided, we can directly use a simulator to obtain all states and observation values of the physical scene. In the context of circuit analysis, we use SPICE (Nagel, 1975) as our simulation language. For other scenarios, there are corresponding choices, such as ANSYS APDL (Kohnke, 1982) in mechanical disciplines and ZPL (Laikin, 2018) in optics domains. Specifically, we develop an expert visual language model to complete this conversion. Since there is no large-scale available dataset or existing model for this task, we devise a data synthesis pipeline to generate abundant physical diagrams and their corresponding simulation languages for our visual language model training.
Solution to Issue in Analysis: Reasoning under the assistance of simulation. Although current MLLMs can perform mathematical reasoning using external tools (Chen et al., 2022; Zhou et al., 2023), recent research found it still challenging to prompt LLMs to write programs for solving scientific problems (Tian et al., 2024). In the benchmark evaluating real-world scientific programming tasks, even the best model achieves an accuracy of less than 10% in completing a main problem.
To address the issue of hallucination when MLLMs perform mathematical derivations and synthesize scientific programs, we delegate the main quantitative reasoning tasks to a domain-specific tool, namely a physical simulator. The simulator comprises domain-specific knowledge and thus is guaranteed to be precise in its output with respect to the given input.
@@ 4.1 Implementation
Synthesis of Training Data for PPM. In the context of circuit diagrams, the diagram layout is defined as the planar grid and components connected between the grid nodes, with the values or labels annotated alongside the component symbols. The grid structures of synthetic data are randomly sampled from a predefined hierarchical distribution, ensuring the diversity of shapes, components, and annotations in the generated circuits.
We use CircuitTikz as our plotting language to draw the circuit diagram using a LaTeX compiler. Since the component annotations in the real-world diagrams can be in a numerical format (e.g. $10\Omega$) or a label format (e.g. $R_{1}$), we generate two types of circuit diagrams to cope with this variation accordingly: (1) Numerical-type circuit, where the value is annotated on the diagram. The PPM is required to infer both the type and value of the components; (2) Label-type circuit, where only the labels of the components are provided in the diagram. The PPM predicts the type and label of the components, with an <Empty> token in the value position.
The physical structure of a circuit diagram can be represented using by a netlist model (Nagel, 1975; Tao et al., 2024), which is a directed graph where each node represents an equipotential point, and each edge represents a circuit component. The SL synthesis step involves writing rules to automatically identify equivalent circuit nodes using basic physical properties and converting grid information into a netlist. We utilize SPICE (Nagel, 1975) as our simulation language for circuit analysis problems. The syntax of SPICE is based on a netlist model, allowing us directly translated the netlist model into SPICE (Nagel, 1975) program at the end of each generation process. Please refer to Appendix B.1 for our design of the hierarchical distribution and a detailed illustration of the synthesis process.
@@
PPM Training. For the training of PPM, we adopt CogVLM-17B (Wang et al., 2023a) as the base model of PPM. The PPM is fine-tuned to generate the SPICE description for given the circuit diagram. For our detailed settings, please refer to Appendix B. Based on our preliminary experiments, the base model is largely unable to accurately perform the conversion task for most circuit diagrams when using prompting methods. Therefore, the training stage is essential for the development of the MAPS pipeline.
Inference. In our main experiments, we use GPT-4V as our MLLM and NgSPICE (Nenzi & Vogt, 2011) as our physical simulator to execute circuit simulation. Given a circuit analysis problem with a diagram and textual description, our framework infers the answer to the problem following the process described in Algorithm 1. For more implementation details, please refer to Appendix A.
@@ 4.2 Evaluation of PPM
Simulation Accuracy ($\text{ACC}_{\text{sim}}$). This metric measures correctness of PPM’s conversion results by comparing the consistency of simulation results between the generated SPICE code and the label code. Although $\text{ACC}_{\text{CQ}}$ is a necessary condition for PPM to be useful in MAPS, in practice, achieving the same simulation results indicates the same physical circuit with high probability. For the specific examples of these metrics, please refer to Appendix B.2.
@@ 4.3 Evaluation of MAPS Framework
We found that our framework and baseline methods all fail at solving problems collected from the Chapter 2 of the textbook, which mainly focuses on the Equivalent Transformation of Resistance and mostly cover the circuits that could not be directly executed in a simulator. When the problem is not simulatable, the MLLM can leverage the additional information from simulation language to reason the final answer. Please refer to our Appendix C.2 for more specific case studies. However, how to improve MLLM’s general scientific reasoning ability through interaction with a physical simulator is still a challenging problem and remains for our future work.
@@ 5.1 Analysis on Inference Phase Design of MAPS
A: We found that when the problem is not simulatable, the simulation language can still be helpful to the final reasoning of framework. The structural information provided by the simulation language significant reduces hallucination of MLLM when understanding the diagram, akin to the role of scene graph in general multi-modal reasoning (Mitra et al., 2024). Appendix C.2 presents a detailed example showing how simulation description in MAPS alleviate the MLLM’s hallucination problem when the physical scene is not simulatable.
@@ 5.2 Analysis of PPM Construction
6 Related Work
@@
7 Discussion & Conclusion
Appendix A Additional Implementation Details of MAPS Inference Phase
A.1 Simulation
@@ A.2 Prompt Templates
Appendix B Additional Implementation Details of PPM Construction Phase
@@ B.2 PPM Training
Appendix C Additional Details of Evaluation on SimpleCircuitEval
C.1 Details of SimpleCircuitEval

### v2

@@ Abstract
Pre-trained on extensive text and image corpora, current Multi-Modal Large Language Models (MLLM) have shown strong capabilities in general visual reasoning tasks. However, their performance is still lacking in physical domains that require understanding diagrams with complex physical structures and quantitative analysis based on multi-modal information. To address this, we develop a new framework, named Multi-Modal Scientific ReAsoning with Physics Perception and Simulation (MAPS) based on an MLLM. MAPS decomposes expert-level multi-modal reasoning task into physical diagram understanding via a Physical Perception Model (PPM) and reasoning with physical knowledge via a simulator. The PPM module is obtained by fine-tuning a visual language model using carefully designed synthetic data with paired physical diagrams and corresponding simulation language descriptions. At the inference stage, MAPS integrates the simulation language description of the input diagram provided by PPM and results obtained through a Chain-of-Simulation process with MLLM to derive the underlying rationale and the final answer. Validated using our collected college-level circuit analysis problems, MAPS significantly improves reasoning accuracy of MLLM and outperforms all existing models. The results confirm MAPS offers a promising direction for enhancing multi-modal scientific reasoning ability of MLLMs. Our code is available at https://github.com/thu-coai/MAPS.
@@ 1 Introduction
To address the aforementioned limitations, we proposed Multi-Modal Scientific ReAsoning with Physics Perception and Simulation (MAPS), a novel framework for solving complex multi-modal reasoning problems in physical disciplines. And in this paper we verified the effectiveness of MAPS in electrical discipline, which typically involves multiple circuit diagrams and is representative of expert-level physical science requiring reasoning on complex physical diagrams. The core idea of MAPS is to decompose expert-level reasoning problems into two sub-tasks: understanding the physical diagram and reasoning based on this comprehension and related physical knowledge. MAPS realizes physical diagram understanding by fine-tuning a visual language model using carefully designed synthetic data, resulting in what we term the Physics Perception Model (PPM). The role of PPM is to translate physical diagrams into simulation language descriptions that can be executed by a simulator. At the inference stage, MAPS integrates the converted simulation language description and their respective simulation results, obtained through a Chain-of-Simulation process, to derive intermediate rationales and ultimately the final answer to the question. Experiment results on college-level circuit analysis problems demonstrate that our framework can successfully address the challenges in complex multi-modal reasoning tasks in physical science. Most importantly, it significantly reduces the occurrence of hallucination when using and solving physical equations. This advancement creates new avenues for precise multi-modal scientific reasoning using MLLMs.
@@ 2 Motivation
Solution to Issue in Perception: Translate physical diagrams into simulation language descriptions. We adopt simulation language for two reasons. First, it describes the physical scene of the diagram using a formal language, enabling the language model to directly access the fundamental structure behind the question. Second, with parameters of physical objects provided, we can directly use a simulator to obtain all states and observation values of the physical scene. In the context of circuit analysis, we use SPICE (Nagel, 1975) as our simulation language. For other scenarios, there are corresponding choices, such as ANSYS APDL (Kohnke, 1982) in mechanical disciplines and ZPL (Laikin, 2018) in optics domains. Specifically, we develop an expert visual language model to complete this conversion. Since there is no large-scale available dataset or existing model for this task, we devise a data synthesis pipeline to generate abundant physical diagrams and their corresponding simulation languages for our visual language model training.
Solution to Issue in Analysis: Reasoning under the assistance of simulation. Although current MLLMs can perform mathematical reasoning using external tools (Chen et al., 2022; Zhou et al., 2023), recent research found it still challenging to prompt LLMs to write programs for solving scientific problems (Tian et al., 2024). In the benchmark evaluating real-world scientific programming tasks, even the best model achieves an accuracy of less than 10% in completing a main problem. To address the issue of hallucination when MLLMs perform mathematical derivations and synthesize scientific programs, we delegate the main quantitative reasoning tasks to a domain-specific tool, namely a physical simulator. The simulator comprises domain-specific knowledge and thus is guaranteed to be precise in its output with respect to the given input.
@@ 4.1 Implementation
Synthesis of Training Data for PPM. In the context of circuit diagrams, the diagram layout is defined as the planar grid and components connected between the grid nodes, with the values or labels annotated alongside the component symbols. The grid structures of synthetic data are randomly sampled from a predefined hierarchical distribution, ensuring the diversity of shapes, components, and annotations in the generated circuits. We use CircuitTikz as our plotting language to draw the circuit diagram using a LaTeX compiler. Since the component annotations in the real-world diagrams can be in a numerical format (e.g. $10\Omega$) or a label format (e.g. $R_{1}$), we generate two types of circuit diagrams to cope with this variation accordingly: (1) Numerical-type circuit, where the value is annotated on the diagram. The PPM is required to infer both the type and value of the components; (2) Label-type circuit, where only the labels of the components are provided in the diagram. The PPM predicts the type and label of the components, with an <Empty> token in the value position.
The physical structure of a circuit diagram can be represented using by a netlist model (Nagel, 1975; Tao et al., 2024), which is a directed graph where each node represents an equipotential point, and each edge represents a circuit component. The SL synthesis step involves writing rules to automatically identify equivalent circuit nodes using basic physical properties and converting grid information into a netlist. We utilize SPICE (Nagel, 1975) as our simulation language for circuit analysis problems. The syntax of SPICE is based on a netlist model, allowing us directly translated the netlist model into SPICE (Nagel, 1975) program at the end of each generation process. Please refer to Appendix C.1 for our design of the hierarchical distribution and a detailed illustration of the synthesis process.
@@
PPM Training. For the training of PPM, we adopt CogVLM-17B (Wang et al., 2023a) as the base model of PPM. The PPM is fine-tuned to generate the SPICE description for given the circuit diagram. For our detailed settings, please refer to Appendix C. Based on our preliminary experiments, the base model is largely unable to accurately perform the conversion task for most circuit diagrams when using prompting methods. Therefore, the training stage is essential for the development of the MAPS pipeline.
Inference. In our main experiments, we use GPT-4V as our MLLM and NgSPICE (Nenzi & Vogt, 2011) as our physical simulator to execute circuit simulation. Given a circuit analysis problem with a diagram and textual description, our framework infers the answer to the problem following the process described in Algorithm 1. For more implementation details, please refer to Appendix B.
@@ 4.2 Evaluation of PPM
Simulation Accuracy ($\text{ACC}_{\text{sim}}$). This metric measures correctness of PPM’s conversion results by comparing the consistency of simulation results between the generated SPICE code and the label code. Although $\text{ACC}_{\text{CQ}}$ is a necessary condition for PPM to be useful in MAPS, in practice, achieving the same simulation results indicates the same physical circuit with high probability. For the specific examples of these metrics, please refer to Appendix C.2.
@@ 4.3 Evaluation of MAPS Framework
We found that our framework and baseline methods all fail at solving problems collected from the Chapter 2 of the textbook, which mainly focuses on the Equivalent Transformation of Resistance and mostly cover the circuits that could not be directly executed in a simulator. When the problem is not simulatable, the MLLM can leverage the additional information from simulation language to reason the final answer. Please refer to our Appendix D.2 for more specific case studies. However, how to improve MLLM’s general scientific reasoning ability through interaction with a physical simulator is still a challenging problem and remains for our future work.
@@ 5.1 Analysis on Inference Phase Design of MAPS
A: We found that when the problem is not simulatable, the simulation language can still be helpful to the final reasoning of framework. The structural information provided by the simulation language significant reduces hallucination of MLLM when understanding the diagram, akin to the role of scene graph in general multi-modal reasoning (Mitra et al., 2024). Appendix D.2 presents a detailed example showing how simulation description in MAPS alleviate the MLLM’s hallucination problem when the physical scene is not simulatable.
@@ 5.2 Analysis of PPM Construction
6 Discussion & Conclusion
7 Acknowledgement
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604) and Tsinghua University Initiative Scientific Research Program.
Appendix A Related Work
@@
Appendix B Additional Implementation Details of MAPS Inference Phase
B.1 Simulation
@@ B.2 Prompt Templates
Appendix C Additional Implementation Details of PPM Construction Phase
@@ C.2 PPM Training
Appendix D Additional Details of Evaluation on SimpleCircuitEval
D.1 Details of SimpleCircuitEval